Obtain the Fourier series for the following periodic function which has a period of 2π: f(x)=x for−π≤x≤π

1
Expert's answer
2022-04-01T13:41:52-0400

"f(x)=x,[-\pi,\pi]"

The function f(x) is not even, so a0=0, an=0.

The numbers a0, an and bn (n=1,2,...) are called the Fourier coefficients of the function f.

The Fourier series of the function f(x) on the interval (0;T) in terms of the sines of multiple arcs is called the series:

"f(x)=\sum b_n*sin(\frac{\pi*n*x}{T})"

"b_n=\frac{2}{T}\displaystyle\int_0^Tf(x)*sin(\frac{\pi*n*x}{T})dx"

For our data:

"b_n=\frac{2}{\pi}\displaystyle\int_0^\pi f(x)*sin(\frac{\pi*n*x}{\pi})dx=\frac{2}{\pi}\displaystyle\int_0^\pi x*sin(n*x)dx="

"\frac{2}{\pi}(-\pi*\frac{cos(\pi*n)}{n}+\frac{sin(\pi*n)}{n^2}-0)=-2*\frac{(-1)^n}{n}"

Finally

"f(x)=0.0+\sum(-2*\frac{(-1)^n}{n})*sin(n*x)"